OpenStudy (anonymous):

Solve 2 log2 2 + 2 log2 6 –log2 3x = 3

OpenStudy (anonymous):

michele

OpenStudy (anonymous):

@Michele_Laino help

OpenStudy (anonymous):

x=16 yes

OpenStudy (michele_laino):

here, if we apply the properties of logarithms, we can rewrite the left side as below: \[\Large \begin{gathered} {\log _2}{2^2} + {\log _2}{6^2} - {\log _2}3x = {\log _2}\left( {\frac{{4 \times 36}}{{3x}}} \right) = \hfill \\ = {\log _2}\left( {\frac{{144}}{{3x}}} \right) \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

more fractions

OpenStudy (michele_laino):

furthermore, we have: \[\Large 3 = {\log _2}8\]

OpenStudy (anonymous):

x=16 not right

OpenStudy (michele_laino):

your equation, can be rewritten as below: \[\Large {\log _2}\left( {\frac{{144}}{{3x}}} \right) = {\log _2}8\]

OpenStudy (michele_laino):

so, using the uniqueness of logarithm we have to solve this equation: \[\Large \frac{{144}}{{3x}} = 8\] what is x?

OpenStudy (anonymous):

oh no i don't know how to do this fraction michele

OpenStudy (michele_laino):

we have to multiply both sides by 3x, so we get: \[144 = 24x\]

OpenStudy (anonymous):

6!

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

good!

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

i got 100% !!!

OpenStudy (michele_laino):

ok! :):)